# CriticBench: Evaluating Large Language Models as Critic

### v3

@@ 3.3 Reference Critique Generation
After collecting instructions and the corresponding responses, we collect reference critiques on these responses to make the subjective evaluation more reliable, with the assistance of GPT-4.
@@
Due to the connection between feedback and correction dimension, GPT-4 is utilized to generate feedback and corrections sequentially (Step 3 (c) and (e) in Fig. 3). We observe that responses in mathematical and coding tasks pose significant challenges for critiques. Thus, ground-truth answers are provided for GPT-4 as references to generate high-quality feedback and corrections. Then, these feedback and corrections are revised by human annotators and used as references in subjective evaluation, respectively.
Comparison-based Feedback
To investigate the impact of the quality discrepancy between two responses’ qualities on $F_{c}$, given responses $(R_{\rm low},R_{\rm med},R_{\rm high})$ of different quality for each instruction $I$, two pairs are created: $(R_{\rm low},R_{\rm high})$, $(R_{\rm med},R_{\rm high})$. Then we prompt GPT-4 to provide preferences and textual analysis on these pairs (Step 3 (f) in Figure 3), followed by human annotators revising these labels and critiques.
@@
Since recent studies Wang et al. (2023c); Zhang et al. (2024) prove that GPT-4 might incorrectly evaluate the quality of critiques, we ask three human experts to directly provide their Likert scores for generated critiques, ranging from $1$ to $7$, respectively. Due to the complexity of the meta-feedback dimension, we do not collect its textual critiques and leave it for future research.
Statistics of Human Annotation
The average revision rates of textual critiques for the feedback, comparison, and correction are $25.22\%$, $34.83\%$, and $48.37\%$, respectively, indicating that GPT-4’s textual feedback is more likely to be accepted by human annotators, followed by the comparison-based feedback. In contrast, GPT-4’s correction might contain more flaws that need to be corrected. Moreover, the average inner correlation score among annotators is $0.79$, indicating solid inner agreement. More details about the statistics of CriticBench and human annotation protocol can be found in Appendix D and E, respectively.
@@ 4.1 Objective Evaluation
aim to evaluate the consistency between generated scores and average human annotation scores. It is worthy noting that LLMs are prompted with the annotated reference critiques on the meta-feedback dimension. The widely-used Spearman correlations Zar (2005) are computed Lan et al. (2020); Tao et al. (2017); Liu et al. (2023c); Fu et al. (2023), which ranges from $-1$ to $1$ (normalize to $(-100,100)$). Higher scores indicate a higher consistency with human judgments.
Comparison-based feedback evaluation
[…]
@@
is only conducted on mathematics and coding tasks since the revision could be easily verified with the ground-truth answers and the test cases. Thus, the objective evaluation metric is implemented as the pass rate: $\frac{N_{\text{Pass}}}{N}$, where $N$ and $N_{\text{Pass}}$ are the number of the total samples and passed samples, respectively.
More details about the objective evaluation in CriticBench are in Appendix G.
@@ 4.2 Subjective Evaluation
The subjective evaluation aims to evaluate the quality of the generated textual critiques. Since responses in mathematics and coding tasks can be verified, we only conduct subjective evaluations on other $5$ tasks for the correction dimension. Following previous works Wang et al. (2023c); Cui et al. (2023); Li et al. (2024), GPT-4 is leveraged to score the generated critiques. Based on our observations, GPT-4 may not analyze all the flaws effectively. Thus, we prompt GPT-4 with annotated critiques as a reference, facilitating a more reliable subjective evaluation. The case study in Appendix H.1 proves the effectiveness of annotated reference critiques in CriticBench, and more detailed proving is left for future work. It is well-known that LLMs like GPT-4 prefer longer generations during their automatic evaluation, i.e., length bias Wang et al. (2023d); Zeng et al. (2024). However, Figure 5 in Appendix F proves that there is no clue that GPT-4, with our concise and precise reference critiques as input, would give higher scores to longer critiques.
The subjective scores range from $1$ to $10$. Scores of $1$ and $10$ denote the worst and best performance, respectively. To effectively measure high-quality critiques, we set the quality score of human-annotated critiques as $8$, and the critiques with scores $>8$ are better than reference critiques.
@@ 4.3 Overall Score
The overall scores of LLMs on these four critique dimensions are calculated to reflect their overall capability. For objective evaluation, the overall scores are the average scores over four critique dimensions, ranging from 00 to $100$. Since Spearman scores range from $-1$ to $1$, they are first normalized into $(0,100)$ before averaging. For subjective evaluation, since meta-feedback does not require textual critiques, the overall scores are set as the average of feedback, comparison-based feedback, and correction scores, ranging from $1$ to $10$.
@@ 5 Evaluation and Analysis
Extensive experimental results for some representative LLMs on the test set are shown in this section, including 1,010 and 1,214 samples for subjective and objective evaluation, respectively. Due to the page limitation, complete evaluation results of all LLMs on the test and dev set for each task and each critique dimension are placed in Appendix I. Except for the overall analysis of closed-source and open-source LLMs in Section 5.2, we also mainly analyze several intriguing phenomena about the relationship between LLM’s critique ability and some critical factors, including the application tasks (Section 5.3), the difficulty of critique dimensions (Section 5.4), the response quality (Section 5.5), and the scaling law (Section 5.6).
@@ 5.1 LLMs to be Evaluated
We extensively evaluate widely used open-source and closed-source LLMs of different sizes on CriticBench to understand the current progress in this field. For closed-source LLMs, we test GPT-4, Claude, Gemini-Pro, PaLM, GPT-3.5-turbo, etc. For open-source LLMs, we test different model series so that we could also study scaling laws, including Mistral Jiang et al. (2024), LLaMA2 Touvron et al. (2023), Baichuan2 Baichuan (2023), Qwen Bai et al. (2023), InternLM2 Team (2023), WizardLM Xu et al. (2024), Vicuna Chiang et al. (2023), Yi, and DeepSeek DeepSeek-AI (2024), etc. Critique-tuned LLMs are also included Cui et al. (2023); Li et al. (2024) to analyze the effectiveness of the critique corpus. Moreover, we show CriticBench can also be used to evaluate reward models that are fundamental component of RLHF Ouyang et al. (2022): (1) UltraRM-13B Cui et al. (2023); (2) Ziya-7B Wang et al. (2022); (3) SteamSHP Ethayarajh et al. (2022).
Model Feedback Correction Comp-Feedback Avg.
Low Med. High Corr. Avg. Low Med. High Avg. EAvg. SAvg. Easy Hard Avg.
Closed-source LLM
GPT-4-turbo 8.39 8.08 7.86 6.07 7.84 7.91 7.54 7.63 7.69 7.86 8.00 8.30 7.99 8.04 7.86
Claude-instant-1 6.49 5.76 5.29 6.17 5.88 7.74 7.69 7.73 7.72
6.67
6.66
6.16 5.66 5.76 6.45
GPT-3.5-turbo 5.80 4.73 4.63 6.04 5.21 7.61 7.34 7.63 7.55
4.95
5.50
5.15 4.84 4.92 5.89
Open-source LLM (Larger than 30B)
Qwen-72B-Chat 6.29 5.28 5.01 5.92 5.57 7.56 7.29 7.51 7.45
6.08
6.18
5.20 5.00 5.02 6.01
DeepSeek-67B-Chat 6.21 5.23 5.35 5.39 5.53 7.48 7.26 7.18 7.30
5.09
6.04
5.08 4.54 4.69 5.84
Mixtral-8x7B-v0.1 5.74 5.14 4.88 5.76 5.31 7.35 7.23 7.40 7.33
5.11
5.93
4.82 4.63 4.62 5.75
Llama2-70B-Chat 3.85 4.22 4.39 4.21 4.12 7.03 7.11 7.17 7.11
5.58
5.54
4.07 3.97 3.95 5.06
Critique-tuned LLM (13B)
Auto-J-13B 4.87 4.38 4.24 3.10 4.21 - - - -
4.98 4.57 4.63 -
UltraCM-13B 4.07 3.88 3.07 4.84 4.12 - - - -
Open-source LLM (13B-20B)
InternLM2-20B-Chat 6.73 5.77 5.68 6.05 6.03 7.35 7.75 7.33 7.48
4.93
5.80
5.40 5.06 5.10 6.20
Qwen-14B-Chat 4.85 4.40 4.56 5.84 4.81 7.28 7.09 7.39 7.25
4.58
4.88
4.00 4.11 3.98 5.35
Llama2-13B-Chat 3.53 3.78 3.94 3.65 3.70 6.92 7.24 7.16 7.11
4.69
4.56
3.52 3.26 3.32 4.71
Open-source LLM (6B-7B)
InternLM2-7B-Chat 5.51 5.11 4.75 5.82 5.20 7.31 7.01 7.19 7.17
4.66
5.50
4.84 4.54 4.62 5.66
Mistral-7B-ins-v0.2 4.90 4.44 4.46 5.36 4.70 7.19 7.23 7.19 7.20
4.82
5.34
4.52 4.24 4.28 5.39
Qwen-7B-Chat 3.59 3.83 4.24 4.96 4.05 6.26 6.34 6.53 6.38
4.10
4.09
3.57 3.48 3.47 4.63
DeepSeek-7B-Chat 3.33 3.22 3.36 4.22 3.44 6.25 5.66 6.26 6.06
4.09
4.23
3.75 3.56 3.60 4.37
Llama2-7B-Chat 3.26 3.65 3.52 3.49 3.44 6.34 6.44 6.02 6.02
4.24
4.00
3.14 3.31 3.21 4.22
Average for Each Quality Granularity
Average 5.14 4.76 4.66 5.11 4.89 7.17 7.08 7.15 7.12
5.16
5.48
4.78 4.55 4.58 5.56
Table 1: Subjective evaluation results (1-10) on the test set of CriticBench. LLM’s evaluation results are shown on samples with different qualities. Light gray denotes the worst and second-worst performance, while dark gray denotes the best and second-best performance. Best average scores are highlighted in bold. Last column (Avg.) contains the overall scores. EAvg and SAvg denote correction with empty and LLM’s self-feedback, respectively. Since partial models are not optimized for correction and comparison dimensions, their results are not recorded.
Model Feedback ($\boldsymbol{-100,100}$) Correction ($\boldsymbol{\%}$) Comp-Feedback ($\boldsymbol{\%}$) Meta-Feedback ($\boldsymbol{-100,100}$) Avg.
Low Med. High Correct Avg. Low Med. High Avg. EAvg. SAvg. Easy Hard Avg. Low Med. High Avg.
Closed-source LLM
GPT-4-turbo 53.64 61.82 49.98 16.84 63.54 66.88 69.48 72.75 69.67 60.67 52.17 63.98 53.03 57.33 59.85 66.49 60.59 62.90 72.55
GPT-3.5-turbo 43.90 36.56 28.67 18.96 51.44 66.56 61.04 63.34 64.00 18.67 21.00 51.08 33.33 40.67 41.47 18.84 25.88 28.71 60.83
Claude-instant-1 24.75 24.67 21.78 34.47 42.78 49.90 46.70 55.36 50.00 24.84 20.17 55.91 37.12 44.89 49.55 35.70 32.42 38.89 58.93
Open-source LLM ($>$ 30B)
DeepSeek-67B 30.51 21.61 11.56 28.06 42.11 57.71 57.67 50.70 55.00 21.00 21.83 52.69 40.53 45.56 26.95 37.11 34.31 33.60 59.60
Qwen-72B-Chat 25.76 23.74 7.59 36.72 42.64 61.15 47.99 58.09 54.67 14.83 7.83 54.30 36.74 44.00 19.14 37.70 25.11 27.86 58.48
Mixtral-8x7B-v0.1 35.46 39.61 14.09 55.51 51.00 52.81 37.08 44.27 43.34 14.34 14.00 47.85 40.91 43.78 18.84 33.08 28.71 26.66 56.49
Llama2-70B-Chat 21.05 25.79 33.10 20.02 32.79 39.69 38.61 47.22 42.34 8.50 5.33 22.58 20.08 21.11 37.66 28.76 19.81 28.32 48.50
Open-source LLM (13B-20B)
InternLM2-20B 46.69 43.53 25.66 19.00 58.61 62.19 37.71 55.48 50.50 5.67 10.33 52.15 39.39 44.67 0.12 -1.43 26.99 3.95 56.61
Llama2-13B-Chat 11.21 17.63 26.22 37.37 30.61 18.23 30.87 22.55 24.67 5.17 4.50 29.03 18.18 22.67 26.64 14.14 50.02 31.02 44.54
Qwen-14B-Chat -10.28 2.08 16.88 45.21 14.32 33.96 45.21 35.19 38.00 8.33 6.67 16.67 15.15 15.78 15.70 4.96 16.00 10.72 41.58
Critique-tuned LLM (13B)
Auto-J-13B 21.16 32.59 32.54 4.11 36.05 - - - - - - 53.23 46.59 49.33 - - - - -
UltraCM-13B -5.54 7.58 29.97 28.70 21.51 - - - - - - 38.17 37.88 38.00 - - - - -
Open-source LLM (6B-7B)
Mistral-7B 33.55 37.16 38.16 25.35 43.66 51.98 38.65 31.62 38.17 11.17 13.00 36.02 21.97 27.88 32.44 31.05 29.84 31.68 50.93
InternLM2-7B 43.23 40.52 13.02 32.95 49.09 49.90 23.37 38.67 36.17 5.67 6.00 33.87 16.67 23.78 8.90 -9.22 20.99 3.17 46.52
DeepSeek-7B -9.39 -0.51 1.7 18.42 8.26 32.08 21.84 46.48 35.00 5.83 8.17 20.43 18.56 19.33 -7.73 -11.77 21.98 -0.31 39.57
Llama2-7B-Chat -0.54 12.03 11.53 36.86 20.81 16.98 23.40 21.11 21.00 5.50 2.83 6.45 4.55 5.33 5.92 -0.07 15.75 5.67 34.89
Qwen-7B-Chat -11.7 -28.88 -23.87 9.97 -8.09 30.52 28.96 34.97 32.33 6.67 1.33 6.99 4.17 5.33 5.48 11.43 24.14 11.73 34.87
Reward Models
UltraRM-13B 47.42 29.33 39.81 18.06 52.33 - - - - - - 65.05 47.35 54.67 - - - - -
Ziya-7B 15.84 11.98 13.42 17.10 25.81 - - - - - - 48.39 34.09 40.00 - - - - -
Average for Each Quality Granularity
Average 21.93 23.10 20.62 26.51 35.75 46.04 40.58 45.19 43.66 14.44 13.01 39.73 29.80 33.90 22.73 19.78 28.84 22.97 50.99
Table 2: Objective evaluation on the test set. Last column (Avg.) contains the overall scores.

### v4

@@ 4.3 Reference Critique Generation and Annotation
After collecting task inputs and responses, four kinds of reference critiques are collected on these responses to make the objective and subjective evaluation in our proposed CriticEval more reliable.
@@
GPT-4-turbo is utilized to generate feedback and corrections sequentially (Step 3 (c) and (e) in Figure 2). The scalar-valued and textual critiques for feedback dimension are collected, denotes as “score” and “text” in Figure 2. Since responses in math reasoning and coding tasks pose significant challenges for critiques during our annotation, ground-truth answers are provided for GPT-4 as references to generate high-quality feedback and corrections. Then, they are carefully reviewed and revised by human annotators.
Comparison
Our empirical finding suggests that pairs of responses pose greater challenges for comparison if they perform similarly. Therefore, two kinds of pairs are first created: $(R_{\rm low},R_{\rm high})$ and $(R_{\rm med},R_{\rm high})$, designated as the easy and hard samples, respectively. Then, GPT-4-turbo is prompted to provide scalar-valued and text critiques on these pairs (Step 3 (f) in Figure 2). These outputs, labeled as “preferred label” and “text” in Figure 2, are then refined by human annotators.
@@
Since GPT-4 has been proven unreliable to evaluate critiques (wang2023shepherd, ; zhang2024selfcontrast, ), three human experts are asked to provide their quality scores for generated critiques (Step 3 (d) in Figure 2).
During human annotation, multiple human experts are asked to follow a rigorous annotation protocol, detailed in Appendix G.1, and the statistics of human annotation for reference critiques are described in Appendix G.5. Besides, several case studies are shown in Appendix H to facilitate a clear understanding of our proposed CriticEval.
@@ 5.1 Objective Evaluation
aim to evaluate the consistency between generated scores and human judgments. This setup facilitates the generation in a chain-of-thought manner, followed by the quality score of the evaluated critiques. For the meta-feedback dimension, LLMs are prompted with annotated reference critiques. The widely-used Spearman correlations (inbook, ) are computed (10.1145/3423168, ; xu-etal-2023-instructscore, ), which ranges from $-1$ to $1$ (normalize to $(-100,100)$). Higher scores indicate a higher consistency with human judgments. The $p$-value of spearman correlation are recorded, and $p<0.05$ is typically considered to be statistically significant 10.1145/3423168 ; Tao2017RUBERAU .
Comparison evaluation
assesses the accuracy of LLM in deciding preferences between two responses. It is well known that current LLMs exhibit significant positional bias (zheng2023judging, ; wang2023large, ; zeng2024evaluating, ), i.e., LLMs tend to prefer responses based on their specific position in the prompt. We implement a rigorous verification process to mitigate the effects of positional bias. Specifically, given responses $R_{a}$ and $R_{b}$ to be compared, we obtain the comparison based on two orders, noted as $F_{c}^{a}=F_{c}(R_{a},R_{b})$ and $F_{c}^{b}=F_{c}(R_{b},R_{a})$. The objective scores are computed by: $s=\frac{1}{N}{\sum_{i=1}^{N}\mathds{1}(L(F_{c}^{a},F_{c}^{b}))}$, where $\mathds{1}\rightarrow\{0,1\}$ is the indicator function. $L(F_{c}^{a},F_{c}^{b})$ is true if and only if $F_{c}^{a}\neq F_{c}^{b}$ and $F_{c}^{a},F_{c}^{b}$ align with ground-truth preference label. $N$ is the number of test samples.
@@
is only conducted on math reasoning and coding tasks since the revision could be easily verified with the ground-truth answers and the test cases (gou2024critic, ; luo2023critique, ). Thus, the objective evaluation metric is implemented as the pass rate: $\frac{N_{\text{Pass}}}{N}$, where $N$ and $N_{\text{Pass}}$ are the number of the total samples and passed samples, respectively.
@@ 5.2 Subjective Evaluation
The subjective evaluation aims to evaluate the quality of the generated textual critiques. Since responses in math reasoning and coding tasks can be verified, we only conduct subjective evaluations on other $5$ tasks for the correction dimension. In our work, GPT-4 evaluates the generated critiques by generating the chain of thought followed by the score, with our human-annotated critiques as references. It is well-known that LLMs prefer longer generations during their automatic evaluation (wang2023how, ; zeng2024evaluating, ). However, Figure 10 in Appendix I proves that there is no clue that GPT-4, with our concise and precise reference critiques as input, would give higher scores to longer critiques. The subjective scores range from $1$ to $10$. Following previous work (liu2023alignbench, ), the human-annotated reference critiques are anchored to $8$, serving as a relative scoring pivot.
@@ 5.3 Overall Score
The overall score of subjective and objective evaluation is computed as averaging on all the critique dimensions, and more details about computing the overall score can be found in Appendix J.
@@ 6 Evaluation and Analysis
Table 2: Subjective and objective evaluation results on the test set of CriticEval. Dark gray and shade gray in this and the following tables highlight the best and worst performance. Objective feedback and meta-feedback scores with $>0.05$ $p$-value are marked with $\dagger$. $\boldsymbol{F_{s}},\boldsymbol{CR},\boldsymbol{F_{c}},\boldsymbol{F_{s}(F_{s})}$ represent feedback, correction, comparison and meta-feedback critique dimensions, respectively. Overall column denotes the overall score over multiple critique dimensions.
Models Subjective Evaluation Objective Evaluation
$\boldsymbol{F_{s}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{c}}$ Overall $\boldsymbol{F_{s}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{c}}$ $\boldsymbol{F_{s}(F_{s})}$ Overall
Closed-source LLM
GPT-4-turbo 7.84 7.69 7.89 7.81 63.54 69.67 57.33 62.90 72.55
GPT-3.5-turbo 5.21 7.55 4.92 5.89 51.44 64.00 40.67 28.71 60.83
Claude-instant-1 5.88 7.72 5.76 6.45 42.78 50.00 44.89 38.89 58.93
Open-source Qwen Series LLMs (bai2023benchmarking, )
Qwen-72B-Chat 5.57 7.45 5.02 6.01 42.64 54.67 44.00 27.86 58.48
Qwen-14B-Chat 4.81 7.25 3.98 5.35 14.32 38.00 15.78 10.72 41.58
Qwen-7B-Chat 4.05 6.38 3.47 4.63 -8.09 32.33 5.33 11.73 34.87
Open-source InternLM2 Series LLMs (2023internlm, )
InternLM2-20B 6.03 7.48 5.10 6.20 58.61 50.50 44.67 3.95 56.61
InternLM2-7B 5.20 7.17 4.62 5.66 49.09 36.17 23.78 3.17 46.52
Open-source Mistral Series LLMs (jiang2024mixtral, )
Mixtral-8x7B 5.31 7.33 4.62 5.75 51.00 43.34 43.78 26.66 56.49
Mistral-7B 4.70 7.20 4.28 5.39 43.66 38.17 27.88 31.68 50.93
Open-source Llama-2 Series LLMs (touvron2023llama, )
Llama2-70B-Chat 4.12 7.11 3.95 5.06 32.79 42.34 21.11 28.32 48.50
Llama2-13B-Chat 3.70 7.11 3.32 4.71 30.61 24.67 22.67 31.02 44.54
Llama2-7B-Chat 3.44 6.02 3.21 4.22 20.81 21.00 5.33 5.67 34.89
The critique abilities of representative LLMs are analyzed in this section, and the overview results are shown in Table 2. Firstly, the reliability of evaluation in CriticEval are proven in Section 6.2 and Section 6.3. Then, overall analysis is described in Section 6.4. Several intriguing phenomena about some critical factors are described, including task types (Section 6.5), the response quality (Section 6.6) and the critique dimensions (Section 6.7). The complete experimental results of all evaluated LLMs on the test/dev set for each task and each critique dimension are placed in Appendix K.
@@ 6.1 LLMs to be Evaluated
35 widely used open-source and closed-source LLMs of different sizes are evaluated on CriticEval, including the instruction-tuned LLMs (qwen, ; deepseek-llm, ; cai2024internlm2, ; openai2023gpt4, ), critique-tuned LLMs that fine-tuned on critique datasets generated by GPT-4 (li2024generative, ; cui2023ultrafeedback, ; jiang2023tigerscore, ), and reward models (cui2023ultrafeedback, ; fengshenbang, ; pmlr-v162-ethayarajh22a, ). Please refer to Appendix L for all evaluated LLMs and the inference details. The prompt templates for LLMs on critique dimensions are shown in Appendix H with score rubrics listed in Figure 17 in Appendix G.2.
6.2 Reliability of Subjective Evaluation in CriticEval
As LLMs are prompted with human-annotated critiques, their performance in meta-feedback could reveal their reliability for evaluating generated textual critiques. As shown in Table 2 ($F_{s}(F_{s})$ column) and Table 3, GPT-4-turbo achieves very high correlations ($62.90$, and $66.18$) with human judgment. Although there is still a gap compared to the average human level ($66.18<79.03$), the strong correlations ensure the reliable evaluation for textual critique ability 10.1145/3423168 . Furthermore, we also conduct the ablation study to prove the contribution of our human-annotated reference critiques. As shown in Table 3, it can be found that all LLMs perform worse when the reference critiques (ref.) are removed (average -$13.36$ performance decrease), proving their significant contribution for reliable subjective evaluation in our proposed CriticEval.
Table 3: Results of meta-feedback dimension in CriticEval dev set. $p$-value $<0.05$.
Models $\boldsymbol{F_{s}(F_{s})}$ $\boldsymbol{F_{s}(F_{s})}$ w/o ref.
GPT-4-turbo 66.18 47.26 (-18.92)
Qwen-1.5-72B 38.97 22.35 (-16.62)
Claude-instant-1 36.88 19.88 (-17.00)
GPT-3.5-turbo 17.28 16.38  (-0.90)
Moreover, except for the feedback critique dimension, we also test the reliability of subjective evaluation on the correction and comparison critique dimensions. Specifically, we ask three human annotators to annotate the quality score of 450 critiques generated by five representative LLMs (GPT-3.5-turbo, Qwen-72B-Chat, InternLM2-20B-Chat, Mistral-7B and ChatGLM3-6B) from 9 tasks in CriticEval, and all the human annotators are guided by the same evaluation protocol in our subjective evaluation.
Table 4: Correlations in $\boldsymbol{CR}$ and $\boldsymbol{F_{c}}$ dimensions. $p$-value $<0.05$.
- $\boldsymbol{CR}$ $\boldsymbol{F_{c}}$
Human Avg. 87.04 76.55
GPT-4 w/ ref. 82.10 70.27
The results are shown in Table 4. It can be found that GPT-4-turbo, with our human-annotated critiques as references, could achieve a very strong correlation with human judgments, close to the average human level. This observation proves the robust and reliable subjective evaluation of the textual critiques in the correction and comparison dimensions. Besides, the correlation scores on the correction and comparison critique dimension are higher than the feedback dimension. This phenomenon suggests that the feedback of the feedback is more challenging than the feedback of correction and comparison.
6.3 More Effective Critiques Consistently Lead to Superior Corrections
Although the reliability of subjective evaluation has been proven in Section 6.2, it is still unknown whether real feedback critique ability of LLMs is consistent with the evaluation results in CriticEval. This section answers this question by demonstrating that better critiques consistently lead to superior corrections. Specifically, two distinct LLMs are investigated: InternLM2-Chat-20B and Llama-2-70B-Chat, which exhibit notable differences in critique abilities in Table 2.
Table 5: The quality of corrections $F_{s}$ increase from top to bottom.
Models Objective Subjective
$\boldsymbol{F_{s}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{s}}$ $\boldsymbol{CR}$
InternLM2+OF 2.24 7.15 5.63 5.71
InternLM2+SF 7.53 10.33 6.85 5.80
InternLM2+HF 8.00 50.50 8.00 7.48
Llama2-70B+SF 2.24 5.33 5.63 5.54
Llama2-70B+OF 5.01 12.47 6.85 6.32
Llama2-70B+HF 8.00 42.34 8.00 7.11
Both models are prompted with three types of feedback $F_{s}$ of varying qualities: (1) Other LLM Feedback (OF), (2) LLM’s Self-generated Feedback (SF), and (3) Human-annotated Feedback (HF) which is anchored to the score of $8$, serving as a relative scoring pivot. The OF for InternLM2-Chat models originates from the Llama-2-70B-Chat, whereas for Llama-2-70B-Chat, OF is sourced from InternLM2-20B-Chat. Therefore, as shown in Table 5, for the InternLM2-20B-Chat model, the critique quality order from high to low is HF$>$SF$>$OF, which reverses to HF$>$OF$>$SF for Llama-2-70B-Chat. Note that quality scores $F_{s}$ of these critiques are evaluated within CriticEval. Subsequently, we prompt these LLMs to correct responses using these three feedback and compute both objective and subjective evaluation scores ($\boldsymbol{CR}$) within CriticEval. As illustrated in Table 5, a clear and consistent trend emerges: better feedback leads to superior corrections. This finding underscores that real critique ability of LLMs aligns closely with the evaluation results in our proposed CriticEval, i.e., critiques of LLMs with higher scores are more accurate and effective for corrections.
In summary, the reliability of evaluation in CriticEval has been well proven. Following sections will describe the overall analysis and relationships between critique ability and several crucial factors.